b) Given certain information, calculate the WACC of the company. tax is 30%

Funding Source

Quantity

Price

Value

Cost

Ordinary Shares

700,000

5.00

18%

Preference Share

250,000

4.50

14%

Retained Earnings

1,000,000

17%

10%, $100 Debentures

11,500

108.00

8%

Bank Overdraft

500,000

10%
Answer #1
1 2 3 4 5 6
Source of income Quantity Price Amount of Investment Weight Cost after tax WACC
2*3 4/TOTAL
Ordinary Share     700,000         5.00                     3,500,000.00 0.48 0.18 0.09
Prefrence Shares     250,000         4.50                     1,125,000.00 0.15 0.14 0.02
Retained Earning             -              -                       1,000,000.00 0.14 0.17 0.02
10% Debenture       11,500      108.00                     1,242,000.00 0.17 0.056 0.01
Bank Overdraft             -              -                         500,000.00 0.07 0.07 0.00
TOTAL                  7,367,000.00 1.00
WACC 0.14
Working Note: WACC 14%
Cost of Debenture   = r(1-tax)(F-P/N) r = Rate of interest
(F+P/N) t = Tax rate applicable
F = Face value of debenture
P= Current Market price of debenture
n= Redemption period of debenture